Circle the equation that are perpendicular.
y=5/4x-6, y=10, -4/5x-y=9, 5x-11y=7, 4y=5x-40

Answer by stanbon(75887)

Circle the equation that are perpendicular.
y=5/4x-6,
slope = 5/4
---------------------
y=10,
slope = 0
---------------------
-4/5x-y=9,
y = (-4/5)x-9
slope = -4/5
---------------------
5x-11y=7,
y = (5/11)x-(7/11)
slope = 5/11
------------------------
4y=5x-40
y = (5/4)x-40
slope = 5/4
---------------------
Lines with slopes of 4/5 and -5/4 are perpendicular
Lines with slopes of 5/4 and -4/5 are perpendicular
=======================================================
Cheers,
Stan H.
==============